OpenStudy (thetruesolition):

f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ please help!!!!!!!!!!!

OpenStudy (loser66):

\[f(\color{red}{n})= f(1)+f(2) +f(\color{red}{n}-1)\] so that \[f(\color{red}{5})= f(1)+f(2) +f(\color{red}{5}-1)\\=f(\color{red}{5})= f(1)+f(2) +f(4)\] you have f(1) =2, f(2)= 5 all you need is f(4) . so find f(4) by exactly the way you find f(5). You will have f(3) there. :) So?? repeat the process to get f(3), plug back to take f(4) , again, plug back to get f(5) Practice to get the concept. :) Although it is a little bit long, but there is no other way to find out. You have to handle it by yourself.